Horizontal Analysis

The comparative accounts payable and long-term debt balances of a company are provided below.

Current Year Previous Year
Accounts payable $52,076 $55,400
Long-term debt 77,066 68,200

Based on this information, what is the amount and percentage of increase or decrease that would be shown in a balance sheet with horizontal analysis? Enter all answers as positive numbers.

Amount of Change Increase/Decrease Percentage
Accounts payable $ %
Long-term debt $ %
Answer #1

Calculate following

Amount of change Increase / Decrease Percentage
Account payable (52076-55400) = -3324 Decrease 3324/55400 = -6%
Long term debt (77066-68200) = 8866 Increase 8866/68200 = 13%